solve the system of equations using the addition method
9x-5y=4
27x-15y=8

Answer by RAY100(1637)

To use the addition method we first must multiply the 1st eqn by 3
.
(1) (9x -5y =4) * 3 = 27x -15y =12
.
(2) 27x -15y =8
.
.
If we subtract (1) from (2), we will get 8=12,,,which is never true
.
Therefore we have no solutions
.
.
another way to look at this is that both eqns are the same except for the y intercepts, therefore they are parallel lines,,,also with no solution
.